OpenStudy (anonymous):

Find the product (2x - 1)^3

OpenStudy (anonymous):

=6x-3

OpenStudy (anonymous):

i dont think that correct

OpenStudy (radar):

(2x-1)(2x-1)(2x-1) or (2x-1)^2(2x-1) (4x^2-4x+1)(2x-1) 8x^3-8x^2+2x-4x^2+4x-1 combing similar terms 8x^3-12x^2+6x-1 If I kept things in order.

OpenStudy (anonymous):

(a-b)^3= a^3 - 3.a^2.b +3.a.b^2 -b^3 now put a=2x and b=1 we will get your answer

OpenStudy (radar):

Ima gonna check that out. (2x)^3-3(2x)^2(1)+3(2x)(1)-1 8x^3-12x^2+6x-1 That too